OpenStudy (abbycross167):

Find the slope-intercept form of the equation that goes through these two points- (1, 20) (8, 4.5) Can someone help me?

OpenStudy (benlindquist):

Do you have a graph

OpenStudy (benlindquist):

it helps

OpenStudy (smartanne):

Do you know the slope equation?

OpenStudy (smartanne):

Just plug in the x and y coordinates in the equation and you will get the slope

OpenStudy (benlindquist):

You need the equation, and a chart if you want to make it easy

OpenStudy (smartanne):

Once you know the slope you can plug in one of the coordinates in the slope intercept equation (y=mx+b) and you will get the y intercept

OpenStudy (yaya090600):

abby? :O

OpenStudy (smartanne):

hope this helps!

OpenStudy (abbycross167):

I do not have a graph... what does the slope equation look like?

OpenStudy (benlindquist):

Hard to do without one

OpenStudy (abbycross167):

oh ok @benlindquist

OpenStudy (anonymous):

We want an equation of the form y=mx+c, so we need to find m (the gradient) and c (the y intercept). Given our two points, we can find the gradient (the change in y divided by the change in x): \[m=\dfrac{20-4.5}{1-8}=\dfrac{15.5}{-7}=-\dfrac{31}{14}\]Now we have m, we can use this and either of our coordinates to find c.\[y=mx+c=-\dfrac{31}{14}x+c\]Let's use (1,20):\[20=-\dfrac{31}{14}\times1+c\]\[c=20+\dfrac{31}{14}=\dfrac{311}{14}\]Giving our answer:\[y=-\dfrac{31}{14} x+\dfrac{311}{14}\]

OpenStudy (smartanne):

\[y ^{2}-y ^{1}\div x ^{2}-x ^{1}\]

OpenStudy (smartanne):

thats the slope equation

OpenStudy (abbycross167):

Ok thank you @Smartanne

OpenStudy (anonymous):

@Smartanne, you should really use subscript for variables otherwise it looks like you're raising it to that power.\[\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (abbycross167):

@tom982 In the answer you gave me those are fractions right?

OpenStudy (anonymous):

Yes, that's in the line where I worked out m, the gradient.

OpenStudy (smartanne):

ok i'll do that next time @tom982